# Context-free graphs and their transition groups

### v2

@@ 9. Free product of context-free inverse graphs
Let $\Theta_{1}=\{(\Gamma^{(i)},\gamma^{i}_{0}),i=1,\ldots,k\}$ be a collection of $A_{1}$-digraphs, and similarly let $\Theta_{2}=\{(\Lambda^{(j)},\lambda^{(j)}_{0}),j=1,\ldots,m\}$ be a collection of $A_{2}$-digraphs, with $A_{1}\cap A_{2}=\emptyset$. For each $i\in[1,k]$, $j\in[1,m]$ we consider the functions $\psi^{1}_{i}:V(\Gamma^{(i)})\setminus\{\gamma_{0}^{i}\}\to\{\lambda^{1}_{0},\ldots,\lambda^{m}_{0}\}$ and $\psi^{2}_{j}:V(\Lambda^{(j)})\setminus\{\lambda_{0}^{j}\}\to\{\gamma^{1}_{0},\ldots,\gamma^{k}_{0}\}$; the gluing map is the collection $\Psi$ of all such functions. We first consider the infinite disjoint union of copies of the graphs $\Theta_{1},\Theta_{2}$ where the set of vertices are words alternating between vertices of the graphs $\Theta_{1}$ and $\Theta_{2}$. More precisely, we consider the $A_{1}\cup A_{2}$-digraph $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ whose set of vertices $\mathcal{V}$ is formed by alternating sequences of vertices of $\Theta_{1},\Theta_{2}$ which are either of the form $\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{\ell}},\mbox{ or }\lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{\ell}}$, with $\gamma^{i_{n}}_{s_{n}}\in V(\Gamma^{(i_{n})}),\lambda^{j_{n}}_{t_{n}}\in V(\Lambda^{(j_{n})})$, $n\in[1,\ell]$ and $\ell\in\mathbb{N}$. The edges are of the form
$\begin{cases}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{h}},a\in A_{2},\mbox{ whenever }\lambda^{j_{\ell}}_{t_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda^{j_{\ell}}_{t_{h}}\in E(\Lambda^{(j_{\ell})})\\ \lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{h}},a\in A_{1},\mbox{ whenever }\gamma^{j_{\ell}}_{s_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma^{j_{\ell}}_{s_{h}}\in E(\Gamma^{(j_{\ell})})\end{cases}$
Roughly speaking we are taking infinitely many copies of the graphs $\Gamma^{(i)},\Lambda^{(j)}$ where we simply rename each vertex by adding a prefix that is a word that is alternating in the sense explained above. Now, the free product $\Theta_{1}\star_{\Psi}\Theta_{2}$ is the $A_{1}\cup A_{2}$-digraph obtained from $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ by identifying the vertices of $\mathcal{V}$ via the transitive closure of the following relation:
@@
Roughly speaking the rooted graph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$ is obtained from the graphs $\Gamma^{(1)}\sqcup\Lambda^{(1)}$ by first identifying $\gamma_{0}^{1}=\lambda_{0}^{1}$, and then by iteratively gluing at a vertex with representative $v^{*}\gamma^{j}_{i}$, with $i\neq 0$ (not a root), the graph $(\Lambda^{(s)},\lambda^{s}_{0})$ with $\lambda^{s}_{0}=\psi^{1}_{j}(\gamma^{j}_{i})$, so that the two vertices $v^{*}\gamma_{i}^{j},v^{*}\lambda_{0}^{s}$ identify. In case the vertex ends with a vertex from the $\Lambda$ graphs such as $v^{*}\lambda^{j}_{i}$, then we glue the graph $(\Gamma^{(s)},\gamma_{0}^{s})$ by making the identification $\psi^{2}_{j}(\lambda^{j}_{i})=\gamma_{0}^{s}$. Let us call a vertex $v^{*}\gamma^{j}_{i}$, where the last vertex belongs to some graph of the collection $\Theta_{1}$, a vertex colored by $1$, while if it ends by a vertex from the collection $\Theta_{2}$ we say that it is colored by $2$. To have an idea of what kind of digraph we may generate see Fig. 2 in which the free product of two rooted graphs $(\Gamma,\gamma_{0}),(\Lambda,\lambda_{0})$ is depicted, where the gluing maps are $\psi^{1}(v)=\lambda_{0}$ for all $v\in\Gamma\setminus\{\gamma_{0}\}$ and $\psi^{2}(v)=\gamma_{0}$ for all $v\in\Lambda\setminus\{\lambda_{0}\}$. We will see later that if $\Gamma,\Lambda$ are two Cayley graphs of two groups $G_{1},G_{2}$, then $(\Gamma,\gamma_{0})\star_{\Psi}(\Lambda,\lambda_{0})$ is the Cayley graph of the free product $G_{1}\star G_{2}$, hence the name. The following proposition is an easy consequence of the definition and the condition $A_{1}\cap A_{2}=\emptyset$.
@@
If $\Theta_{1}$ ($\Theta_{2}$) is a family of deterministic (complete) $A_{1}$-digraphs ($A_{2}$-digraphs), then $\Theta_{1}\star_{\Psi}\Theta_{2}$ is a deterministic (complete) $A_{1}\cup A_{2}$-digraph. In particular, if $\Theta_{1}$ ($\Theta_{2}$) is a family of inverse $A_{1}$-graphs ($A_{2}$-graphs), then the rooted graphs $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$ is an inverse $A_{1}\cup A_{2}$-graph.
We now make some geometric considerations on this construction. For simplicity of notation let $\alpha=\gamma_{0}^{1}=\lambda_{0}^{1}$ and put $\Omega=\Theta_{1}\star_{\Psi}\Theta_{2}$. Now, consider a geodesic $\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}$ connecting $\alpha$ to some vertex $v^{*}\gamma_{i}^{j}$ of $(\Omega,\alpha)$, colored by $1$ with $i\neq 0$. Since $i\neq 0$, $w$ ends with a word colored by $1$, thus it may be uniquely factor as an alternating product $w=z_{1}x_{1}z_{2}x_{2}\ldots z_{k}x_{k}$ with $z_{t}\in A_{2}$ and $x_{t}\in A_{1}$. Let $\Lambda^{(s)}$ be the digraph of $\Theta_{2}$ with $\lambda_{0}^{s}=\psi^{1}_{j}(\gamma_{i}^{j})$ so that in $\Omega$, the vertex $v^{*}\gamma_{i}^{j}$ is identified with $v^{*}\lambda_{0}^{s}$. Note that we are gluing at $v^{*}\gamma_{i}^{j}$ the rooted $A_{1}\cup A_{2}$-digraph $\Omega^{+}(v^{*}\gamma_{i}^{j})$ obtained from $\Lambda^{(s)}$ in which we iteratively glue all the digraphs from the collection $\Theta_{1},\Theta_{2}$ according to the gluing maps $\Psi$. More formally, this subgraph is isomorphic to the $A_{1}\cup A_{2}$-digraph $\Lambda^{(s)}(\uparrow_{\Psi})$ which is the connected component containing $\lambda_{0}^{s}$ of the digraph obtained from $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ by identifying the vertices of $\mathcal{V}$ via the equivalence relation generated by the relation
@@
Note that $\Lambda^{(s)}(\uparrow_{\Psi})\simeq\Omega^{+}(v^{*}\gamma_{i}^{j})$ via the map on the set of vertices defined by $\phi(\lambda_{0}^{s})=v^{*}\gamma_{i}^{j}=v^{*}\lambda_{0}^{s}$ and $\phi(\lambda_{0}^{s}u^{*})=v^{*}\lambda_{0}^{s}u^{*}$. The digraph $\Omega^{+}(v^{*}\gamma_{i}^{j})$ is called the out-cone of $\Omega$ at $v^{*}\gamma_{i}^{j}$. Moreover, since we are gluing a copy of $\Lambda^{(s)}(\uparrow_{\Psi})$ by identifying the vertex $\lambda_{0}^{s}$ with $v^{*}\gamma_{i}^{j}$, this means that all the geodesics $\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}$ of $\Omega$ passing through $v^{*}\gamma_{i}^{j}$ with $w^{\prime}=u_{1}u_{2}$, for some $u_{1}\in A_{2}^{+}$, have the property that $v^{*}\gamma_{i}^{j}u^{*}$ belongs to the out-cone $\Omega^{+}(v^{*}\gamma_{i}^{j})$, see Fig. 3 for a graphical representation. We record this remark in the following lemma.
@@
$V(\Omega^{+}(v^{*}\gamma_{i}^{j}))=\{v^{*}\gamma_{i}^{j}\}\cup\{v^{*}\gamma_{i}^{j}u^{*}:\mbox{ there is a geodesic }\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}\mbox{ with }w^{\prime}\in A_{2}^{+}(A_{1}\cup A_{2})^{*}\}$
$V(\Omega^{+}(v^{*}\lambda_{i}^{j}))=\{v^{*}\lambda_{i}^{j}\}\cup\{v^{*}\lambda_{i}^{j}u^{*}:\mbox{ there is a geodesic }\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\lambda_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\lambda_{i}^{j}u^{*}\mbox{ with }w^{\prime}\in A_{1}^{+}(A_{1}\cup A_{2})^{*}\}$
@@
Figure 3. In the dashed region the out-cone $\Omega^{+}(v^{*}\gamma_{i}^{j})$, in red the geodesic $\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}$.
@@
With the above notation, for a given vertex $v^{*}\gamma_{i}^{j}\lambda^{s}_{t}$ of $\Omega$ colored by $2$, the end-cone $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ is isomorphic to the $A_{1}\cup A_{2}$-digraph obtained by taking the disjoint union

### v3

@@ 9. Free product of context-free graphs
[…]
$\begin{cases}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{\ell}}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{h}},a\in A_{2},\mbox{ whenever }\lambda^{j_{\ell}}_{t_{\ell}}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\lambda^{j_{\ell}}_{t_{h}}\in E(\Lambda^{(j_{\ell})})\\ \lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{\ell}}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{h}},a\in A_{1},\mbox{ whenever }\gamma^{j_{\ell}}_{s_{\ell}}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$a$}}}{{\relbar\joinrel\longrightarrow}}}\gamma^{j_{\ell}}_{s_{h}}\in E(\Gamma^{(j_{\ell})})\end{cases}$
Roughly speaking we are taking infinitely many copies of the graphs $\Gamma^{(i)},\Lambda^{(j)}$ where we simply rename each vertex by adding a prefix that is a word that is alternating in the sense explained above. Now, the free product $\Theta_{1}\star_{\Psi}\Theta_{2}$ is the $A_{1}\cup A_{2}$-graph obtained from $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ by identifying the vertices of $\mathcal{V}$ via the transitive closure of the following relation:
@@
Roughly speaking the rooted graph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$ is obtained from the graphs $\Gamma^{(1)}\sqcup\Lambda^{(1)}$ by first identifying $\gamma_{0}^{1}=\lambda_{0}^{1}$, and then by iteratively gluing at a vertex with representative $v^{*}\gamma^{j}_{i}$, with $i\neq 0$ (not a root), the graph $(\Lambda^{(s)},\lambda^{s}_{0})$ with $\lambda^{s}_{0}=\psi^{1}_{j}(\gamma^{j}_{i})$, so that the two vertices $v^{*}\gamma_{i}^{j},v^{*}\lambda_{0}^{s}$ identify. In case the vertex ends with a vertex from the $\Lambda$ graphs such as $v^{*}\lambda^{j}_{i}$, then we glue the graph $(\Gamma^{(s)},\gamma_{0}^{s})$ by making the identification $\psi^{2}_{j}(\lambda^{j}_{i})=\gamma_{0}^{s}$. Let us call a vertex $v^{*}\gamma^{j}_{i}$, where the last vertex belongs to some graph of the collection $\Theta_{1}$, a vertex colored by $1$, while if it ends by a vertex from the collection $\Theta_{2}$ we say that it is colored by $2$. To have an idea of what kind of graph we may generate see Fig. 2 in which the free product of two rooted graphs $(\Gamma,\gamma_{0}),(\Lambda,\lambda_{0})$ is depicted, where the gluing maps are $\psi^{1}(v)=\lambda_{0}$ for all $v\in\Gamma\setminus\{\gamma_{0}\}$ and $\psi^{2}(v)=\gamma_{0}$ for all $v\in\Lambda\setminus\{\lambda_{0}\}$. We will see later that if $\Gamma,\Lambda$ are two Cayley graphs of two groups $G_{1},G_{2}$, then $(\Gamma,\gamma_{0})\star_{\Psi}(\Lambda,\lambda_{0})$ is the Cayley graph of the free product $G_{1}\star G_{2}$, hence the name. The following proposition is an easy consequence of the definition and the condition $A_{1}\cap A_{2}=\emptyset$.
@@
If $\Theta_{1}$ ($\Theta_{2}$) is a family of deterministic (complete) $A_{1}$-graphs ($A_{2}$-graphs), then $\Theta_{1}\star_{\Psi}\Theta_{2}$ is a deterministic (complete) $A_{1}\cup A_{2}$-graph. In particular, if $\Theta_{1}$ ($\Theta_{2}$) is a family of $A_{1}$-graphs ($A_{2}$-graphs), then the rooted graphs $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$ is a $A_{1}\cup A_{2}$-graph.
We now make some geometric considerations on this construction. For simplicity of notation let $\alpha=\gamma_{0}^{1}=\lambda_{0}^{1}$ and put $\Omega=\Theta_{1}\star_{\Psi}\Theta_{2}$. Now, consider a geodesic $\alpha\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}$ connecting $\alpha$ to some vertex $v^{*}\gamma_{i}^{j}$ of $(\Omega,\alpha)$, colored by $1$ with $i\neq 0$. Since $i\neq 0$, $w$ ends with a word colored by $1$, thus it may be uniquely factor as an alternating product $w=z_{1}x_{1}z_{2}x_{2}\ldots z_{k}x_{k}$ with $z_{t}\in A_{2}$ and $x_{t}\in A_{1}$. Let $\Lambda^{(s)}$ be the graph of $\Theta_{2}$ with $\lambda_{0}^{s}=\psi^{1}_{j}(\gamma_{i}^{j})$ so that in $\Omega$, the vertex $v^{*}\gamma_{i}^{j}$ is identified with $v^{*}\lambda_{0}^{s}$. Note that we are gluing at $v^{*}\gamma_{i}^{j}$ the rooted $A_{1}\cup A_{2}$-graph $\Omega^{+}(v^{*}\gamma_{i}^{j})$ obtained from $\Lambda^{(s)}$ in which we iteratively glue all the graphs from the collection $\Theta_{1},\Theta_{2}$ according to the gluing maps $\Psi$. More formally, this subgraph is isomorphic to the $A_{1}\cup A_{2}$-graph $\Lambda^{(s)}(\uparrow_{\Psi})$ which is the connected component containing $\lambda_{0}^{s}$ of the graph obtained from $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ by identifying the vertices of $\mathcal{V}$ via the equivalence relation generated by the relation
@@
Note that $\Lambda^{(s)}(\uparrow_{\Psi})\simeq\Omega^{+}(v^{*}\gamma_{i}^{j})$ via the map on the set of vertices defined by $\phi(\lambda_{0}^{s})=v^{*}\gamma_{i}^{j}=v^{*}\lambda_{0}^{s}$ and $\phi(\lambda_{0}^{s}u^{*})=v^{*}\lambda_{0}^{s}u^{*}$. The graph $\Omega^{+}(v^{*}\gamma_{i}^{j})$ is called the out-cone of $\Omega$ at $v^{*}\gamma_{i}^{j}$. Moreover, since we are gluing a copy of $\Lambda^{(s)}(\uparrow_{\Psi})$ by identifying the vertex $\lambda_{0}^{s}$ with $v^{*}\gamma_{i}^{j}$, this means that all the geodesics $\alpha\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w^{\prime}$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}$ of $\Omega$ passing through $v^{*}\gamma_{i}^{j}$ with $w^{\prime}=u_{1}u_{2}$, for some $u_{1}\in A_{2}^{+}$, have the property that $v^{*}\gamma_{i}^{j}u^{*}$ belongs to the out-cone $\Omega^{+}(v^{*}\gamma_{i}^{j})$, see Fig. 3 for a graphical representation. We record this remark in the following lemma.
@@
$V(\Omega^{+}(v^{*}\gamma_{i}^{j}))=\{v^{*}\gamma_{i}^{j}\}\cup\{v^{*}\gamma_{i}^{j}u^{*}:\mbox{ there is a geodesic }\alpha\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w^{\prime}$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}\mbox{ with }w^{\prime}\in A_{2}^{+}(A_{1}\cup A_{2})^{*}\}$
$V(\Omega^{+}(v^{*}\lambda_{i}^{j}))=\{v^{*}\lambda_{i}^{j}\}\cup\{v^{*}\lambda_{i}^{j}u^{*}:\mbox{ there is a geodesic }\alpha\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\lambda_{i}^{j}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w^{\prime}$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\lambda_{i}^{j}u^{*}\mbox{ with }w^{\prime}\in A_{1}^{+}(A_{1}\cup A_{2})^{*}\}$
@@
Figure 3. In the dashed region the out-cone $\Omega^{+}(v^{*}\gamma_{i}^{j})$, in red the geodesic $\alpha\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$w^{\prime}$}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}$.
@@
With the above notation, for a given vertex $v^{*}\gamma_{i}^{j}\lambda^{s}_{t}$ of $\Omega$ colored by $2$, the end-cone $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ is isomorphic to the $A_{1}\cup A_{2}$-graph obtained by taking the disjoint union